Determine the value of n which make the following statment true:
(1/2) nP7= n-2P5

Answer by stanbon(75887)

Determine the value of n which make the following statment true:
(1/2) nP7= n-2P5
(1/2)n!/(n-7)! = (n-2)!/(n-7)!
---
The denominators are equal so equate the numerators:
(1/2)n! = (n-2)!
Divide both sides by (n-2)!
---------
(1/2)n(n-1) = 1
n(n-1) = 2
n^2-n-2 = 0
(n-2)(n+1) = 0
Positive solution:
n = 2
=========================
Hmmmm? That answer does not make sense because
you cannot have 2P7 in the original equation.
=========================
Cheers,
Stan H.